# Hidden Minima in Two-Layer ReLU Networks

### v3

@@ Abstract.
We consider the optimization problem arising from fitting two-layer ReLU networks with $d$ inputs under the square loss, where labels are generated by a target network. Two infinite families of spurious minima have recently been identified: one whose loss vanishes as $d\to\infty$, and another whose loss remains bounded away from zero. The latter are nevertheless avoided by vanilla SGD, and thus hidden, motivating the search for analytic properties distinguishing the two types. Perhaps surprisingly, the Hessian spectra of hidden and non-hidden minima agree up to terms of order $O(d^{-1/2})$, providing limited explanatory power. Consequently, our analysis of hidden minima proceeds instead via curves along which the loss is minimized or maximized. The main result is that arcs emanating from hidden minima differ, characteristically, by their structure and symmetry, precisely on account of the $O(d^{-1/2})$-eigenvalue terms absent from previous analyses.
@@ 1. Introduction
where $\sigma(x)\coloneqq\max\{0,x\}$ is the ReLU function acting entrywise and $M(k,d)$ denotes the space of $k\times d$ matrices. To isolate the study of optimization-related obstructions on account of nonconvexity from those pertaining to the expressive power of two-layer networks, data has been often assumed to be fully realizable. This was further motivated by hardness results indicating a strict barrier inherent to the explanatory power of distribution-free approaches operating in complete generality [13, 16, 43]. For the squared loss, the resulting highly nonconvex expected loss is
@@
In [2], the theory of distributions plays a central role in addressing the nonsmooth nature of ReLU networks when computing expectations and their derivatives for a general class of inputs, including the normal distribution. This viewpoint provides a unified and rigorous framework for differentiating with respect to network parameters, including biases. For the sample loss, tools from variational analysis complement this framework and play a key role in capturing the pertinent nonsmooth optimization structure.
@@ 4. Numerical results: bounding the distance to the nearest critical point
In practice, we construct tangency arcs for $\mho_{C}({\mathcal{L}})$, $C\in M(d,d)$ a critical point, using a Lagrangian function encoding a norm constraint, $Q(W,\eta,r)={\mathcal{L}}(W)+\eta(\|W-C\|^{2}-r^{2})$. The (augmented) tangency arcs $(W(r),\eta(r))$ are computed by solving $DQ=0$ using Newton-Raphson method and small increments of $r$, typically 1e-3, until hitting a singular Jacobian or until $r_{\text{max}}$ has been reached. The former case indicates the presence of a critical point or an arc bifurcating (both occur!). The latter case is taken to indicate our (very) finite approximation for an arc continuing indefinitely. The limit $r_{\text{max}}$ has been chosen so as to be an order of magnitude larger than the typical length of arcs terminating at ‘finite’ time. Following Lemma 1, arcs are initialized by setting $W(0)\coloneqq C+r_{\text{min}}B$ where $B$ denotes an eigenvector and $r_{\text{min}}=\text{1e-7}$. Repeating this process for different minima and ranges of eigenvalues, we found that arcs corresponding to eigenvectors associated to small eigenvalues tend to be finite and generally terminate earlier than those corresponding to large eigenvalues. In Table 4, we report the radius of arcs corresponding to minimal eigenvalues. The numerical estimates are consistent with Theorem 2:
@@
5. Concluding remarks
@@
The focus in this paper has been on $m(r)$ and $M(r)$, giving respectively the minimal and maximal value of the loss function in the vicinity of critical points. The functions are studied as a means of identifying analytic properties differentiating hidden minima, type I, from ones detected by standard gradient-based methods, type II. We prove general results on tangency arcs realizing $m(r)$ and $M(r)$ showing how pure group representation-theoretic considerations yield a precise description of the admissible types of such arcs. The general results used for the loss function $\mathcal{L}$} reveal that tangency arcs of type I differ from type II by their structure and symmetry, provably requiring a greater extent of symmetry breaking to have curves realizing $m(r)$ on account of $O(d^{-1/2})$-terms of the Hessian spectrum of type I and II which are otherwise identical. The theoretical results are derived using methods developed for o-minimal structures which imply in particular a certain topological regularity required for the numerical work presented. In addition to confirming the structure and symmetry type predicted by our theoretical results, the construction of tangency arcs gives an effective means for estimating the distance from a given critical point to adjacent ones.
@@
Critical points of $G$-invariant functions may or may not be symmetric. In [1], it is shown, however, that when certain conditions apply, critical points connected to symmetric ones by tangency arcs are symmetry breaking. Once a lower bound on the isotropy of critical points has been obtained, the effects of increasing the number of neurons on spurious minima and, crucially, on the possible emergence of descent directions can be analyzed, e.g., [8]. A systematic study of these phenomena for loss landscapes arising from deep architectures is undertaken in [2], where the theory of distributions provides a coherent and rigorous framework for addressing the nonsmooth behavior of ReLU networks when computing expectations and their derivatives, across all parameters, including biases. For the sample loss, complementary tools from variational analysis join this framework and further refine the description of the underlying nonsmooth optimization structure.

### v4

@@ Abstract.
We consider the optimization problem associated with training two-layer ReLU networks with $d$ inputs under the squared loss, where the labels are generated by a target network. Recent work has identified two distinct classes of infinite families of minima: one whose training loss vanishes in the high-dimensional limit, and another whose loss remains bounded away from zero. The latter family is empirically avoided by stochastic gradient descent, hence hidden, motivating the search for analytic criteria that distinguish hidden from non-hidden minima. A key challenge is that prior analyses have shown the Hessian spectra at hidden and non-hidden minima to coincide up to terms of order $O(d^{-1/2})$, seemingly limiting the discriminative power of spectral methods. We therefore take a different route, studying instead certain curves along which the loss is locally minimized. Our main result shows that arcs emanating from hidden minima exhibit distinctive structural and symmetry properties, arising precisely from $\Omega(d^{-1/2})$ eigenvalue contributions that are absent from earlier analyses.
@@ 1. Introduction
where $\sigma(t)\coloneqq\max\{0,t\}$ is the ReLU function acting entrywise and $M(k,d)$ denotes the space of $k\times d$ matrices. To isolate the study of optimization-related obstructions on account of nonconvexity from those pertaining to the expressive power of two-layer networks, data has been often assumed to be fully realizable. This was further motivated by hardness results indicating a strict barrier inherent to the explanatory power of distribution-free approaches operating in complete generality [13, 16, 43]. For the squared loss, the resulting highly nonconvex expected loss is
@@
In [3], the theory of distributions is adopted as a central analytical framework for the rigorous treatment of the nonsmoothness inherent in ReLU networks. This approach applies to a broad class of data distributions given by Radon measures, that is, finite Borel measures of total mass one, and in particular includes the normal distribution as is the case in the present work. It furnishes a unified and mathematically precise foundation for differentiation with respect to network parameters, encompassing, in particular, bias terms (see the concluding remark in Section 5 below). Within this framework, techniques from geometric measure theory provide a refined structural characterization of the gradient, which is viewed as a vector-valued function of bounded variation, as well as of the Hessian, now represented in the distributional sense as a matrix-valued Radon measure.
@@ 4. Numerical results: bounding the distance to the nearest critical point
In practice, we construct tangency arcs for $\mho_{C}({\mathcal{L}})$, $C\in M(d,d)$ a critical point, using a Lagrangian function encoding a norm constraint, $Q(W,\eta,r)={\mathcal{L}}(W)+\eta(\|W-C\|^{2}-r^{2})$. The (augmented) tangency arcs $(W(r),\eta(r))$ are computed by solving $DQ=0$ using Newton-Raphson method and small increments of $r$, typically 1e-3, until hitting a singular Jacobian or until $r_{\text{max}}$ has been reached. The former case indicates the presence of a critical point or an arc bifurcating (both occur). The latter case is taken to indicate our (very) finite approximation for an arc continuing indefinitely. The limit $r_{\text{max}}$ has been chosen so as to be an order of magnitude larger than the typical length of arcs terminating at ‘finite’ time. Following Lemma 1, arcs are initialized by setting $W(0)\coloneqq C+r_{\text{min}}B$ where $B$ denotes an eigenvector and $r_{\text{min}}=10^{-7}$. Repeating this process for different minima and ranges of eigenvalues, we found that arcs corresponding to eigenvectors associated to small eigenvalues tend to be finite and generally terminate earlier than those corresponding to large eigenvalues. In Table 4, we report the radius of arcs corresponding to minimal eigenvalues. The numerical estimates are consistent with Theorem 2:
@@
5. Concluding remarks and future work
@@
The focus in this paper has been on $m(r)$ and $M(r)$, giving respectively the minimal and maximal value of the loss function in the vicinity of critical points. The functions are studied as a means of identifying analytic properties differentiating hidden minima, type I, from ones detected by standard gradient-based methods, type II. We prove general results on tangency arcs realizing $m(r)$ and $M(r)$ showing how pure group representation-theoretic considerations yield a precise description of the admissible types of such arcs. The general results used for the loss function $\mathcal{L}$} reveal that tangency arcs of type I differ from type II by their structure and symmetry, provably requiring a greater extent of symmetry breaking to have curves realizing $m(r)$ on account of $O(d^{-1/2})$-terms of the Hessian spectrum of type I and II which are otherwise identical. The theoretical results are derived using methods developed for o-minimal structures which imply in particular a certain topological regularity required for the numerical work presented. In addition to confirming the structure and symmetry type predicted by our theoretical results, the construction of tangency arcs provides an effective means for studying critical points adjacent to a given one.
Bias terms
A natural question arising from the analysis of minima in this work is how the conclusions drawn in the unbiased setting are affected by the introduction of bias terms. The biased analogue of the kernel appearing in the unbiased analysis is as follows. For $\mathbf{w},\mathbf{v}\in\mathbb{R}^{d}$ and $a,b\in\mathbb{R}$, define
$k(\mathbf{w},a,\mathbf{v},b):=\mathbb{E}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},I_{d})}\!\left[\sigma(\langle\mathbf{w},\mathbf{x}\rangle+a)\,\sigma(\langle\mathbf{v},\mathbf{x}\rangle+b)\right].$
Assume $\|w\|>0$ and $\|v\|>0$ (the edge cases follow similarly). Set $\rho=\langle\mathbf{w},\mathbf{v}\rangle/(\|\mathbf{w}\|\|\mathbf{v}\|)$, $\alpha=-a/\|\mathbf{w}\|$, $\beta=-b/\|\mathbf{v}\|$, and $s=\sqrt{1-\rho^{2}}$. Let $\phi$ and $\Phi$ denote the standard normal density and distribution functions, and let $\Phi_{2}(\cdot,\cdot;\rho)$ denote the bivariate normal distribution function with correlation $\rho$. Define also the bivariate normal density
$\phi_{2}(u,v;\rho)=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\exp\!\left(-\frac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\right).$
Then the kernel admits the explicit expression
$k(\mathbf{w},a,\mathbf{v},b)={}$ $\ \|\mathbf{w}\|\,\|\mathbf{v}\|\Bigg(\rho\Big[1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)\Big]+(1-\rho^{2})\,\phi_{2}(\alpha,\beta;\rho)\Bigg)$
$\quad+\ \|\mathbf{w}\|\,b\;\phi(\alpha)\Big(1-\Phi\!\big(\tfrac{\beta-\rho\alpha}{s}\big)\Big)+\ \|\mathbf{v}\|\,a\;\phi(\beta)\Big(1-\Phi\!\big(\tfrac{\alpha-\rho\beta}{s}\big)\Big)$
$\quad+\ ab\Big[1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)\Big].$
Remark 2 (Owen’s $T$-function).
The bivariate normal distribution function $\Phi_{2}$ can also be expressed using Owen’s $T$-function, which yields equivalent one-dimensional integral representations that are often employed for numerical evaluation.
First- and second-order derivatives of the kernel likewise admit explicit expressions. For example, the gradient with respect to $\mathbf{w}$ takes the form
$\nabla_{\mathbf{w}}k(\mathbf{w},a,\mathbf{v},b)=c_{1}\,\mathbf{w}+c_{2}\,\mathbf{v},$
where
$\Delta$ $\coloneqq\langle\mathbf{w},\mathbf{w}\rangle\langle\mathbf{v},\mathbf{v}\rangle-\langle\mathbf{w},\mathbf{v}\rangle^{2}=\|\mathbf{w}\|^{2}\|\mathbf{v}\|^{2}(1-\rho^{2}),$
$c_{1}$ $=\frac{\langle\mathbf{v},\mathbf{v}\rangle\,\Gamma_{1}-\langle\mathbf{w},\mathbf{v}\rangle\,\Gamma_{2}}{\Delta},\qquad c_{2}=\frac{-\langle\mathbf{w},\mathbf{v}\rangle\,\Gamma_{1}+\langle\mathbf{w},\mathbf{w}\rangle\,\Gamma_{2}}{\Delta}.$
$\Gamma_{1}$ $=\|\mathbf{w}\|\Bigg(\|\mathbf{v}\|\Big(\rho\big[1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)\big]+(1-\rho^{2})\phi_{2}(\alpha,\beta;\rho)\Big)$
$+b\,\phi(\alpha)\Big(1-\Phi\!\big(\tfrac{\beta-\rho\alpha}{s}\big)\Big)\Bigg),$
$\Gamma_{2}$ $=\|\mathbf{v}\|\Bigg(\|\mathbf{v}\|\Big(1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)+(1-\rho^{2})\,\beta\,\phi(\beta)\Big(1-\Phi\!\big(\tfrac{\alpha-\rho\beta}{s}\big)\Big)$
$+\rho(1-\rho^{2})\phi_{2}(\alpha,\beta;\rho)\Big)+b\,\phi(\beta)\Big(1-\Phi\!\big(\tfrac{\alpha-\rho\beta}{s}\big)\Big)\Bigg).$
A detailed and systematic analysis of the structural and symmetry properties, together with precise asymptotic characterizations of the eigenspectrum, extending the methodology developed in the present work, and a comparative study of the resulting classes of global and spurious minima associated with the biased kernel, is carried out in [2].
Symmetry breaking and deep architectures
@@
Critical points of $G$-invariant functions may or may not be symmetric. In [1], it is shown, however, that when certain conditions apply, critical points connected to symmetric ones by tangency arcs are symmetry breaking. Once a lower bound on the isotropy of critical points has been obtained, the effects of increasing the number of neurons on spurious minima and, crucially, on the possible emergence of descent directions can be analyzed, e.g., [9].
A rigorous and systematic analysis of these phenomena in loss landscapes arising from the training of deep neural architectures is developed in [3]. In that work, the theory of distributions is adopted as a comprehensive analytical framework that permits a precise treatment of the nonsmoothness of ReLU networks. Within this framework, tools from geometric measure theory are employed to obtain a refined characterization of the gradient, interpreted as a vector-valued function of bounded variation, and of the Hessian, which is accordingly represented by Radon measures.